OpenStudy (brunerc1141):

find the slope and y intercept of the equation 4x + 5y = -9

OpenStudy (mjdennis):

If you convert it to slope-intercept form, you will be able to read the answer right from the new form. Slope-intercept form means we need to get from Ax + By = C to y = mx +b To do this, start with your equation 4x + 5y = -9

OpenStudy (mjdennis):

Thanks, @Mehek14 ! Of course, @brunerc1141 has "left the building". I'm going home for the night also. No one is moving very fast on this problem.

OpenStudy (arnavguddu):

express in intercept form (x/A) + (y/B) = 1 given 4x + 5y = -9 so (4x/-9) + (5y/-9) = 1 (x/(-9/4)) + (y/(-9/5)) = 1 So intercept => A = -9/4, B=-9/5 So coordinates are (A, 0) and (0,B) Slope = (0-B)/(A-0) = -B/A put the values and solve them